# The shrinking target and recurrence problem for non-autonomous systems

### v1

@@ 3.1. Proof of Theorem D
Let $(X,\mathcal{B},\mu,T_{{\mathbb{N}}},d)$ be as in Theorem D. While the overall structure of the proof of Theorem D mirrors [18, Theorem 1.3] (which proves the result in the autonomous case) several estimates must be re-established in the broader, non-autonomous context. We actually will prove the following proposition, from which Theorem D easily follows. Throughout, the constants $c_{1},c_{2},\delta$ are the constants arising from the Ahlfors regularity of $\mu$– see (1.5).
@@ 4.1. Background on one-component inner functions
The following lemma is part of a broader equivalences given in [19, Theorem 7.2 (d)].
@@
For finite degree maps, every regular value has exactly $\deg(f)$ preimages. This cannot be directly extended to infinite degree maps without clarifying if the degree of the map is countable or not. In fact, centred, one-component inner functions are essentially countable to one, meaning that for $\lambda$ almost every $z\in\partial{\mathbb{D}}$, $\{f^{-1}(z)\}$ is at most countable [19, Corollary 3.3].
@@
By the Schwarz lemma, $f^{-1}A_{\rho}\subset A_{\rho}$, and by [19, Lemma 7.1], the map $f:{\mathbb{D}}\to{\mathbb{D}}$ is a covering map on this annulus. Hence, every branch of $f^{-1}$ is well defined in $A_{\rho}$. Furthermore, as $f^{-n}A_{\rho}\subset A_{\rho}$, the same is true for $f^{-n}$. ∎
@@ 4.2. Markov Partition for one-component inner functions
In [19, §8], Ivrii-Urbaǹski construct a Markov partition for infinite degree one-component inner functions which generalises the above method for finite Blaschke products. This construction is more delicate because now $\Sigma\neq\emptyset$. Therefore, $\Sigma$ already partitions $\partial{\mathbb{D}}$ into arcs $\{I_{i}\}$. By Lemma 4.2, as $f$ is one-component, its associated exceptional set $\Sigma$ coincides with the singular set $E(f)$ and so is closed. We can decompose the unit circle into open arcs $\partial{\mathbb{D}}\setminus\Sigma=\{I_{i}\}_{i=1}^{\infty}$. By Remark Remark, we also know that all points in $\partial{\mathbb{D}}$ are regular. We first must show that there exists a fixed point $x\in\partial{\mathbb{D}}$. We then will use to generate the arcs $U_{i}$ in (4.3) as before. To this end, pick an arbitrary point $p$ and consider $f^{-1}(p)$. This divides each $I_{i}$ into countably many arcs (as $f$ is essentially countable to one, and acts as a covering map on $I_{i}$ by the discussion surrounding Lemma 4.3) each of which is mapped by $f$ to $\partial{\mathbb{D}}\setminus\{p\}$. This implies that each subdivided arc of $I_{i}$ (which has end points in $f^{-1}(p)$) must contain a fixed point. Therefore, it is possible to find $x$. As before, we use $x$ to define the sets $\{U_{i}\}$ (of which there will be countably many) that generate our partition.
@@ 4.3. Proof of Theorem F
We will split the proof up into two cases based on the degree of $f$ being finite or infinite. In the finite case, the zero-one criterion follows directly from a theorem of He-Liao [16, Theorem 1.4], which applies to piecewise expanding maps. In the one dimensional setting these are defined as follows:
@@
Note that this definition means the map $f$ has to have a finite Markov partition. We state can now state the one dimensional version of [16, Theorem 1.4].
@@
With Lemma 4.15 in hand, it remains to show that positive measure for $\mathcal{R}(f,\{r_{n}\})$ implies full measure. Let $L=\mathcal{R}(f,\{r_{n}\})$ and $L_{m}=\{z\in\partial{\mathbb{D}}:f^{n}(z)\in B(x,r_{n+m})\text{ \ for i.m \ }n\in{\mathbb{N}}\}$. Notice that $f^{-m}(L_{m})=\{z\in\partial{\mathbb{D}}:f^{n+m}(z)\in B(x,r_{n+m})\text{ \ for i.m \ }n\in{\mathbb{N}}\}=L$. As $n$ was arbitrary, by Proposition 2.8, we conclude that the recurrence set has full measure. ∎

### v2

@@ 3.1. Proof of Theorem D
Let $(X,\mathcal{B},\mu,T_{{\mathbb{N}}},d)$ be as in Theorem D. While the overall structure of the proof of Theorem D mirrors [17, Theorem 1.3] (which proves the result in the autonomous case) several estimates must be re-established in the broader, non-autonomous context. We actually will prove the following proposition, from which Theorem D easily follows. Throughout, the constants $c_{1},c_{2},\delta$ are the constants arising from the Ahlfors regularity of $\mu$– see (1.5).
@@ 4.1. Background on one-component inner functions
The following lemma is part of a broader equivalences given in [18, Theorem 7.2 (d)].
@@
For finite degree maps, every regular value has exactly $\deg(f)$ preimages. This cannot be directly extended to infinite degree maps without clarifying if the degree of the map is countable or not. In fact, centred, one-component inner functions are essentially countable to one, meaning that for $\lambda$ almost every $z\in\partial{\mathbb{D}}$, $\{f^{-1}(z)\}$ is at most countable [18, Corollary 3.3].
@@
By the Schwarz lemma, $f^{-1}A_{\rho}\subset A_{\rho}$, and by [18, Lemma 7.1], the map $f:{\mathbb{D}}\to{\mathbb{D}}$ is a covering map on this annulus. Hence, every branch of $f^{-1}$ is well defined in $A_{\rho}$. Furthermore, as $f^{-n}A_{\rho}\subset A_{\rho}$, the same is true for $f^{-n}$. ∎
@@ 4.2. Markov Partition for one-component inner functions
In [18, §8], Ivrii-Urbaǹski construct a Markov partition for infinite degree one-component inner functions which generalises the above method for finite Blaschke products. This construction is more delicate because now $\Sigma\neq\emptyset$. Therefore, $\Sigma$ already partitions $\partial{\mathbb{D}}$ into arcs $\{I_{i}\}$. By Lemma 4.2, as $f$ is one-component, its associated exceptional set $\Sigma$ coincides with the singular set $E(f)$ and so is closed. We can decompose the unit circle into open arcs $\partial{\mathbb{D}}\setminus\Sigma=\{I_{i}\}_{i=1}^{\infty}$. By Remark Remark, we also know that all points in $\partial{\mathbb{D}}$ are regular. We first must show that there exists a fixed point $x\in\partial{\mathbb{D}}$. We then will use to generate the arcs $U_{i}$ in (4.2) as before. To this end, pick an arbitrary point $p$ and consider $f^{-1}(p)$. This divides each $I_{i}$ into countably many arcs (as $f$ is essentially countable to one, and acts as a covering map on $I_{i}$ by the discussion surrounding Lemma 4.3) each of which is mapped by $f$ to $\partial{\mathbb{D}}\setminus\{p\}$. This implies that each subdivided arc of $I_{i}$ (which has end points in $f^{-1}(p)$) must contain a fixed point. Therefore, it is possible to find $x$. As before, we use $x$ to define the sets $\{U_{i}\}$ (of which there will be countably many) that generate our partition.
@@ 4.3. Proof of Theorem F
We will split the proof up into two cases based on the degree of $f$ being finite or infinite. In the finite case, the zero-one criterion follows directly from a theorem of He-Liao [15, Theorem 1.4], which applies to piecewise expanding maps. In the one dimensional setting these are defined as follows:
@@
Note that this definition means the map $f$ has to have a finite Markov partition. We state can now state the one dimensional version of [15, Theorem 1.4].
@@
With Lemma 4.15 in hand, it remains to show that positive measure for $\mathcal{R}(f,\{r_{n}\})$ implies full measure. Let $L=\mathcal{R}(f,\{r_{n}\})$ and $L_{m}=\{z\in\partial{\mathbb{D}}:f^{n}(z)\in B(z,r_{n+m})\text{ \ for i.m \ }n\in{\mathbb{N}}\}$. Notice that $f^{-m}(L_{m})=\{z\in\partial{\mathbb{D}}:f^{n+m}(z)\in B(z,r_{n+m})\text{ \ for i.m \ }n\in{\mathbb{N}}\}=L$. As $n$ was arbitrary, by Proposition 2.8, we conclude that the recurrence set has full measure. ∎